molly has $12.30 in coins. she has four more nickels than dimes. she has 6 fewer quarters than dimes. how many quarters does she have?
Answer by clarinetgirlM(1)

molly has $12.30 in coins. she has four more nickels than dimes. she has 6 fewer quarters than dimes. how many quarters does she have?
a nickel= 0.05, a dime= 0.10, a quarter= 0.25
x = dimes
(x+4) = nickels
(x-6)= quarters
.05(x+4) + .25(x-6) + 0.10(x) = 12.30
.05x + 0.2 + 0.25x - 1.5 + 0.10x = 12.30
.40x - 1.3 = 12.30
.40x = 13.6
x = 34 ( the amount of dimes)
34 x (.10)= $3.40 (is how much of the $12.30 the dimes make up)
For nickels
(x+4)
(34 + 4)= 38 x (.05)= $1.90 ( how much of the $12.30 the nickels make up)
For quarters
(x-6)
(34-6) = 28 x (.25) = $7.00 ( how much of the $12.30 the quarters make up)
And just to check the answers
$3.40 + $1.90 + $7.00 = $12.30
yea so 28 quarters for the answer
ok i hope this helps xD